Suppose that a particle moves along a line with position function s(t) =2t^2 +3t+1 where s is in meters and t is in seconds.


a. What is its initial position?

b. Where is it located after t = 2 seconds?

c. At what time is the particle at position s = 6?




1
Expert's answer
2022-04-27T16:33:18-0400

a.


"s(0)=s_0=2(0)^2 +3(0)+1=1(m)"

b.


"s(2)=2(2)^2 +3(2)+1=15(m)"

15 meters from its initial position.


c.


"s(t)=6 m, t\ge0"

"2t^2 +3t+1=6"

"2t^2 +3t-5=0"

"(t-1)(2t+5)=0"

"t=1\ or\ t=-2.5"

Since "t\ge 0," we take "t=1 s."

The particle will be at position s = 6 after 1 second.